# MRGen: Segmentation Data Engine For Underrepresented MRI Modalities

### v2

@@ 4.2 Quantitative Results
Segmentation. As depicted in Table 3, we draw the following three observations: (i) significant discrepancies between different modalities challenge the generalization ability of nnUNet and UMamba models trained solely on source-domain data, leading to notably lower average DSC scores; (ii) DualNorm and segmentation models trained with data translated by CycleGAN and UNSB, achieve slight or moderate improvements in average DSC scores, but consistently underperform across most scenarios; (iii) conversely, our MRGen produces high-quality target-domain samples for training segmentation models, thus achieving the highest DSC score in 8 out of 10 experiments, significantly outperforming others. Notably, MRGen consistently improves performance of nnUNet and UMamba, demonstrating the adaptability and robustness of its synthetic data across various segmentation architectures. More comparisons will be included in the Sec. D.2.
@@ 4.3 Qualitative Results
Figure 5: Qualitative Results of Segmentation towards Unannotated Modalities. Significant imaging differences between source-domain ($\mathcal{D}_{s}$) and target-domain ($\mathcal{D}_{t}$) make segmentation on target domains ($\mathcal{D}_{t}$) extremely challenging. We visualize liver, right kidney, left kidney and spleen in the first two rows, prostate in the third row, and pancreas in the fourth row with distinct colors.
Segmentation. As illustrated in Figure 5, despite significant appearance variations among distinct modalities, MRGen substantially improves segmentation accuracy across all organs with high-quality synthetic data. However, DualNorm, and segmentation models trained on data derived from CycleGAN and UNSB, yield unsatisfactory results.
@@ 4.4 Ablation Studies
To validate the effectiveness of our strategies and modules, we conduct comprehensive ablation experiments on both generation and downstream segmentation task, as follows.
@@
As presented in Table 4, we can draw the following observations: (i) while SDM pretrained on natural images performs poorly on MRI, finetuning on MRGen-DB yields substantial improvement in both reconstruction and generation; (ii) MRGen, with higher latent dimensions (16 versus 4 of SDM) and the BiomedCLIP text encoder, achieves significantly better performance; and (iii) our templated text prompts further enable MRGen to better distinguish distinct modalities, regions, and organs, leading to superior synthesis quality. To this end, we employ high-capacity autoencoders trained on MRGen-DB, text encoder pretrained on biomedical data, and clinically relevant templated text prompts to ensure accurate and realistic MRI synthesis.
@@
Segmentation. We train nnUNet with data generated by MRGen under various training and inference settings. As shown in Table 5, we have the following two observations: (i) MRGen boosts segmentation performance with synthetic data, even without incorporating target-domain unannotated data during training, demonstrating its strong generalization capability to underrepresented and annotation-scarce modalities; and (ii) the inclusion of target-domain unannotated images and the autofilter pipeline further improve performance by mitigating overfitting and selecting high-quality samples aligned with mask conditions.
@@ Appendix A Preliminaries on Diffusion Models
Here, $\mu_{\theta}$ represents the predicted mean of Gaussian distribution, derived from the estimated noise $\boldsymbol{\epsilon}_{\theta}$ as:
@@ Appendix B Details of MRGen-DB & Synthetic Data
In this section, we provide additional details about our collected and curated MRGen-DB dataset. In Sec. B.1, we elaborate on the implementation details of the automatic annotation pipeline; and in Sec. B.2, we present more comprehensive dataset statistics. Additionally, in Sec. B.3, we provide statistics on the MRGen-synthesized data used for downstream segmentation models training under each experimental setting.
@@ B.1 Automatic Annotations
We employ an automated annotation pipeline to annotate our MRGen-DB dataset, ensuring that the templated text prompts contain sufficient and comprehensive clinically relevant information to distinguish distinct modalities, regions, and organs. This process primarily consists of two precise and controllable components: human body region classification and modality explanation, which will be detailed as follows.
Region classification. Considering the wide range and variability of abdominal imaging, we adopt the off-the-shelf BiomedCLIP [62] image encoder to encode all 2D slices and the BiomedCLIP text encoder to encode predefined text descriptions of six abdominal regions. Based on the cosine similarity between the image and text embeddings, the 2D slices are classified into these six categories, including Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region. For text encoding, we use a templated text prompt as input:
@@ B.3 Synthetic Data Statistics
This section presents the statistics of target-domain training samples synthesized by MRGen across various experimental settings. Concretely, we use mask annotations from the entire source-domain dataset (including both training and test sets) as input conditions to generate target-domain images, forming image-mask training pairs. Exceptions include: (i) for the MSD-Prostate [2] dataset, where images of T2 and ADC modalities have already been registered, we restrict inputs to the source-domain training set to prevent data leakage; and (ii) for dataset pairs with CHAOS-MRI-T1 [26] as the target domain, each source-domain mask is used twice to synthesize both T1-InPhase and T1-OutofPhase data. This setup is consistent across all baselines. Additionally, with our proposed autofilter pipeline, MRGen generates 20 candidate images per mask and selects the top two that meet the predefined thresholds. If no samples satisfy the thresholds, all thresholds will be relaxed by 0.10, and the sample of the highest quality is chosen, ensuring full exploitation of source-domain masks. Otherwise, all low-quality generated samples are discarded to avoid noisy data. Finally, the statistics of target-domain training pairs generated by MRGen for each experimental setting are summarized in Table 7.
@@ C.1 Preprocessing & Augmentation
Data augmentation. During autoencoder training, we apply random data augmentations to images with a 20% probability. These augmentations include horizontal flipping, vertical flipping, and rotations of 90, 180, 270. In contrast, no data augmentations are applied during the training of the diffusion UNet and mask condition controller. We do not explicitly adopt data balancing, as we empirically find that it does not lead to significant performance changes. For downstream segmentation models, we adhere to the default data augmentation strategies provided by nnUNet [23] and UMamba [40].
@@ C.3 Baselines
Additionally, we consider other methods aimed at addressing our focused challenge, i.e., segmenting MR images of underrepresented modalities lacking mask annotations. These include few-shot learning approaches, general-purpose segmentation models, and methods incorporating oracle inputs as performance references. Notably, these approaches, to varying degrees, rely on manually annotated target-domain segmentation masks or external datasets. Thus, they should be regarded as references only rather than fair comparisons with the aforementioned methods and our MRGen.
@@ Appendix D More Experiments
In this section, we present additional experimental results to demonstrate the superiority of our proposed data engine. First, in Sec. D.1, we showcase quantitative and qualitative results of in-domain generation. Next, in Sec. D.2, we present quantitative comparisons with more baselines, further confirming the effectiveness and necessity of our proposed data engine. Finally, in Sec. D.3, we provide extra qualitative results to validate the accuracy and flexibility of the generated outputs.
@@ D.1 In-domain Generation
Our proposed data engine not only synthesizes images for target modalities lacking mask annotations but also maintains controllable generation capabilities within the source domains. Moreover, as presented in Table 8, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real manually-annotated data. This offers a feasible solution to address concerns about medical data privacy.
@@ D.2 More Quantitative Results
D.3 More Qualitative Results
@@
Image generation. As depicted in Figure 9, we present additional visualizations of controllable generation on target modalities lacking mask annotations. These results demonstrate that the proposed data engine can effectively generate high-quality samples based on masks across various datasets and modalities, facilitating the training of downstream segmentation models towards these challenging scenarios.
@@ E.1 Limitations
False-negative samples. Another challenge arises from the varying number of organs to be segmented on each slice. For instance, one slice may contain the liver, kidneys, and spleen, while another may include only the liver and spleen. This variability causes MRGen to occasionally generate additional segmentation targets not specified in the mask condition. For example, as illustrated in Figure 8 (b), kidneys are unexpectedly synthesized by MRGen, despite not being included in the mask conditions, leading to false negatives during the training of downstream segmentation networks. A feasible solution is to design a more comprehensive and robust data filtering pipeline to filter these false-negative samples. Alternatively, simple manual selection can serve as a quick and effective method to remove samples that do not meet the requirements.
@@ E.2 Future Works
Due to limited computational resources, we validate our proposed data engine on 2D slices in this paper, with trained segmentation models able to process 3D volumes slice-by-slice. However, we believe our idea can be seamlessly extended to 3D volume generation. With more computing in the future, we aim to develop a 3D version model for the community, further advancing cross-modality segmentation performance. Moreover, to address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained, accurate, and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust and comprehensive data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
Figure 9: More Qualitative Results of Controllable Generation. We present images from source domains $\mathcal{D}_{s}$ and target domains $\mathcal{D}_{t}$ for reference. Here, specific organs are contoured with colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver in LiQA and CHAOS-MRI datasets.

### v3

@@ 4.2 Quantitative Results
Segmentation. As depicted in Table 3, we draw the following observations: (i) notable discrepancies among different modalities challenge the generalization ability of nnUNet and UMamba trained solely on source-domain data, leading to lower average DSC scores; (ii) DualNorm and segmentation models trained with data translated by CycleGAN and UNSB, achieve slight or moderate improvements, but consistently underperform across most scenarios; (iii) conversely, our MRGen produces high-quality target-domain samples for training segmentation models, thus achieving the best DSC score in 8 out of 10 experiments, significantly outperforming others. Notably, MRGen consistently improves the performance of nnUNet and UMamba, exhibiting the robustness and adaptability of its synthetic data across various segmentation architectures. More comparisons are included in Sec. D.2 of the Appendix.
@@ 4.3 Qualitative Results
Figure 5: Qualitative Results of Segmentation towards Unannotated Modalities. Significant imaging differences between source-domain ($\mathcal{D}_{s}$) and target-domain ($\mathcal{D}_{t}$) make segmentation on target domains ($\mathcal{D}_{t}$) extremely challenging. We visualize liver, right kidney, left kidney, and spleen in the first two rows, prostate in the third row, and pancreas in the fourth row with distinct colors.
Segmentation. As illustrated in Figure 5, despite significant appearance variations among distinct modalities, training on synthetic data from MRGen substantially improves segmentation accuracy across all organs. However, DualNorm, and segmentation models trained on data derived from CycleGAN and UNSB, yield unsatisfactory results.
@@ 4.4 Ablation Studies
To validate the effectiveness of our strategies and modules, we conduct comprehensive ablation experiments on both generation and downstream segmentation tasks, as follows.
@@
As presented in Table 4, we can draw the following observations: (i) while SDM pretrained on natural images performs poorly on MRI, finetuning on MRGen-DB yields substantial improvement in both reconstruction and generation; (ii) MRGen, with higher latent dimensions (16 versus 4 of SDM) and the BiomedCLIP text encoder, achieves significantly better performance; and (iii) our templated text prompts further enable MRGen to better distinguish distinct modalities, regions, and organs, leading to superior synthesis quality. To this end, we employ high-capacity autoencoders trained on MRGen-DB, a text encoder pretrained on biomedical data, and clinically relevant templated text prompts to ensure accurate and realistic MRI synthesis.
@@
Segmentation. We train nnUNet with data generated by MRGen under various training and inference settings. As shown in Table 5, we have the following two observations: (i) MRGen boosts segmentation performance with synthetic data, even without training on MR images from target domain, demonstrating its strong generalization capability to underrepresented and annotation-scarce modalities; and (ii) the inclusion of target-domain MR images (without annotations) and the autofilter pipeline further improve performance by mitigating overfitting and selecting high-quality samples aligned with mask conditions.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Key R & D Program of China (No. 2022ZD0161400).
@@ Appendix A Preliminaries on Diffusion Models
Here, $\mu_{\theta}$ represents the predicted mean of the Gaussian distribution, derived from the estimated noise $\boldsymbol{\epsilon}_{\theta}$ as:
@@ Appendix B Details of MRGen-DB & Synthetic Data
This section provides additional details about our curated MRGen-DB dataset. In Sec. B.1, we elaborate on the implementation details of the automatic annotation pipeline; and in Sec. B.2, we present more comprehensive data statistics. Moreover, in Sec. B.3, we provide statistics on the MRGen-synthesized data used for downstream segmentation models training.
@@ B.1 Automatic Annotations
We employ an automated annotation pipeline to annotate our MRGen-DB dataset, ensuring that the templated text prompts contain sufficient and clinically relevant information to distinguish distinct modalities, regions, and organs. This process primarily consists of two precise and controllable components: human body region classification and modality explanation, which will be detailed as follows.
Region classification. Considering the wide range and variability of abdominal imaging, we adopt the off-the-shelf BiomedCLIP [63] image encoder to encode all 2D slices, and the BiomedCLIP text encoder to encode predefined text descriptions of six abdominal regions. Based on the cosine similarity between the image and text embeddings, the 2D slices are classified into one of the six categories, including Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region. For text encoding, we use a templated text prompt as input:
@@ B.3 Synthetic Data Statistics
This section presents the statistics of target-domain training samples synthesized by MRGen across various experimental settings, as presented in Table 7. Concretely, we use mask annotations from the entire source-domain dataset (including both training and test sets) as input conditions to generate target-domain images, forming image-mask training pairs. Exceptions include: (i) for the MSD-Prostate [3] dataset, where images of T2 and ADC modalities have already been registered, we restrict inputs to the source-domain training set to prevent data leakage; and (ii) for dataset pairs with CHAOS-MRI-T1 [27] as the target domain, each source-domain mask is used twice to synthesize both T1-InPhase and T1-OutofPhase data. This setup is consistent across all baselines. Additionally, with our proposed autofilter pipeline, MRGen generates 20 candidate images per mask and selects the top two that meet the predefined thresholds. If no samples satisfy the thresholds, all thresholds will be relaxed by 0.10, and the sample of the highest quality is chosen, ensuring full exploitation of source-domain masks. Otherwise, all low-quality generated samples are discarded to avoid noisy data.
@@ C.1 Preprocessing & Augmentation
Data augmentation. During autoencoder training, we apply random data augmentations to images with a 20% probability. These augmentations include horizontal flipping, vertical flipping, and rotations of 90, 180, 270. In contrast, no data augmentations are applied during the training of the diffusion UNet and mask condition controller. Our preliminary experiments show that MRGen remains robust to uneven data distribution; we therefore do not explicitly adopt data balancing in training. For segmentation models, we adhere to the default data augmentation strategies provided by nnUNet [23] and UMamba [41].
@@ C.3 Baselines
Additionally, we consider other methods aimed at addressing our focused challenge, i.e., segmenting MR images of underrepresented modalities lacking mask annotations. These include few-shot learning approaches, general-purpose segmentation models, and methods incorporating oracle inputs as performance references. Notably, these approaches, to varying degrees, rely on manually annotated target-domain segmentation masks or external datasets. Thus, they should be regarded as references only, rather than fair comparisons with the aforementioned methods and our MRGen.
@@ Appendix D More Experiments
In this section, we present additional experimental results to demonstrate the superiority of our proposed data engine. First, in Sec. D.1, we showcase quantitative and qualitative results of in-domain generation. Next, in Sec. D.2, we present quantitative comparisons with more baselines, further confirming the effectiveness and necessity of our proposed data engine. Then, in Sec. D.3, we present extra promising application prospects (cross-protocol generation) on paired CT and MRI datasets. Finally, in Sec. D.4, we provide extra qualitative results to validate the accuracy and flexibility of the generated outputs.
@@ D.1 In-domain Generation
Our proposed data engine not only synthesizes images for target modalities lacking mask annotations but also maintains controllable generation capabilities within the source domains. Moreover, as presented in Table 8, downstream segmentation models trained exclusively on synthetic source-domain data can achieve performance comparable to those trained on real, manually-annotated data. This offers a feasible solution to address concerns about medical data privacy.
Figure 7: Qualitative Results of In-domain Generation.
@@ D.2 More Quantitative Results
Source Domain Target Domain DualNorm nnUNet UMamba
$\mathcal{D}_{s}$ CycleGAN MRGen $\mathcal{D}_{s}$ CycleGAN MRGen
AMOS22 (CT) CHAOS (T2) 19.78 0.11 6.75 22.50 0.05 8.06 26.73
AMOS22 (CT) CHAOS (T1) 16.09 8.88 52.49 56.23 3.19 43.21 60.53
MSD-Liver (CT) CHAOS (T2) 1.58 3.12 10.14 38.67 1.65 11.06 40.93
Table 10: Quantitative Results (DSC score) on Cross-protocol settings (from CT to MRI).
D.3 Extension to More Modalities
Considering that evaluation on truly rare modalities is difficult due to limited ground truth annotations, we simulate such scenarios by restricting models from accessing target-modality labels in our experiments. Here, we also explore cross-modality synthesis (from CT to MRI) with AMOS22 [25], MSD-Liver [3], and CHAOS-MRI [27] datasets to further demonstrate MRGen’s potential for broader cross-protocol generation, as depicted in Table 10.
D.4 More Qualitative Results
@@
Image generation. We present extra visualizations of controllable generation on target modalities lacking mask annotations in Figure 10, which demonstrate that our MRGen can effectively generate high-quality samples based on masks across various datasets and modalities, facilitating the training of downstream segmentation models towards these challenging scenarios.
@@ E.1 Limitations
False-negative samples. Another challenge arises from the varying number of organs on each slice. For instance, one slice may contain liver, kidneys, and spleen, while another may include only liver and spleen. This variability causes MRGen to occasionally generate targets not specified in the mask condition. As depicted in Figure 8 (b), kidneys are unexpectedly synthesized, despite not being included in the mask conditions, leading to false negatives during the training of downstream segmentation networks. A feasible solution is to design a more robust data filtering pipeline to filter false-negative samples, and simple manual selection can also serve as a quick and effective method to remove the non-compliant samples.
Figure 9: The validation of 3D extension feasibility of MRGen on $256\times 256\times 16$ volumes.
@@ E.2 Future Works
Due to limited computational resources, we validate our data engine on 2D slices, with trained segmentation models able to process 3D volumes slice-by-slice. However, our idea can be seamlessly extended to 3D volume generation with more computing in the future to further advance cross-modality segmentation performance. Here, we provide a preliminary validation on $256\times 256\times 16$ volumes, as depicted in Figure 9. While the results are not fully optimized due to limited computations, they already demonstrate promising inter-slice consistency, indicating the feasibility of extending MRGen to 3D synthesis.
Moreover, to address the aforementioned limitations of MRGen, we propose several directions for future improvement: (i) Constructing more comprehensive and richly annotated datasets, such as incorporating more annotated MRI data, to enhance the model’s ability to effectively utilize mask conditions; (ii) Designing finer-grained and efficient generative model architectures to improve generation efficiency and accuracy, particularly for small-volume organs; and (iii) Developing a more robust data filtering pipeline to reliably select high-quality samples that meet the requirements of downstream tasks.
Figure 10: More Qualitative Results of Controllable Generation. We present images from source domains ($\mathcal{D}_{s}$) and target domains ($\mathcal{D}_{t}$) for reference. Here, specific organs are contoured with colors: prostate in MSD-Prostate and PROMISE12 datasets, and pancreas in PanSeg dataset, and liver in LiQA and CHAOS-MRI datasets.